June bought a 2 liter cola. she drank 1/3 of a liter on saturday, 1/4 on sunday, 3/8 of a liter on monday. What part of the 2 liter cola does she have left?
Answer by ewatrrr(24785)


 
Hi,
2 - 1/3 - 1/4 - 3/8 = 2 - 8/24 - 6/24 - 9/24 = 2 - 23/24 = = 1 1/24 liter